The length of the rectangle is 4 less than three times the width. If the perimeter is 88cm, what are the dimensions of the rectangle?
Answer by mananth(16946)

let width be x
length will be 3x-4
..
Perimeter = 2*(L+W)
2*(x+3x-4)=88
2*(4x-4)=88
8x-8=88
8x=96
divide by 8
x= 12 cm the width
length = 3x-4 = 36-4=32.
CHECK
2*32+2*12= 88